OpenStudy (anonymous):

Please help! Perform the indicated operations in the following problem; leave the answer in the rectangular form a + b i. Simplify: 7i^2+8i^3+9i^4+10i^5 a. 16+2i b. 2+2i c. 16+18i d. 2+18i

OpenStudy (anonymous):

i^2=-1,i^3=-i,i^4=1,i^5=i Therefor,-7-8i+9+10i=2+2i

OpenStudy (dumbcow):

just use fact that i^2 = -1

OpenStudy (anonymous):

Thanks guys :)